# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ A.2 Helper Lemmas
If Algorithm 1 executes Algorithm 2, then $G$ contains no non-isolating 2-vertex cuts, self loops, parallel edges, irrelevant edges, or $\alpha$-contractable subgraphs with at most $\frac{4}{\varepsilon}$ vertices.
@@
Let $G$ be the graph when Algorithm 1 executes Algorithm 2, $\{u,v,w\}$ be a large 3-vertex cut of $G$, $(V_{1},V_{2})$ be partition of $V\setminus\{u,v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$, and $H$ be a 2EC spanning subgraph of $G$. Let $G_{i}=G[V_{i}\cup\{u,v,w\}]$ and $H_{i}=E(G_{i})\cap H$ for $i\in\{1,2\}$. Then, $H_{1}$ and $H_{2}$ are of a type in $\mathcal{T}$ with respect to  $\{u,v,w\}$ and $G_{1}$ and $G_{2}$, respectively.
@@
First, assume that $H_{i}$ is a single connected component. In this case $H_{i}^{\prime}$ is a tree. If $H_{i}^{\prime}$ is composed of a single node, $H_{i}$ is of type $\mathsf{A}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}$ that are on any simple path between $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$. If $H_{i}^{\prime}$ contains an edge $e\in E(H_{i}^{\prime})\setminus F$, then $e$ cannot belong to any cycle of $H$, because it does not belong to a cycle in $H_{i}$ and is not on a simple path between any two vertices of the cut $\{u,v,w\}$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C1}$ if $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$ are distinct, and of type $\mathsf{B1}$ otherwise.
Second, assume that $H_{i}$ is composed of two connected components. Let $H_{i}^{\prime}(v)$ and $H_{i}^{\prime}(w)$ denote the two trees in $H_{i}^{\prime}$ such that w.l.o.g. $u$ and $v$ are in $H_{i}^{\prime}(v)$ and $w$ is in $H_{i}^{\prime}(w)$. First note that $H_{i}^{\prime}(w)$ must be a single node. Otherwise, there is an edge $e$ in $H_{i}^{\prime}(w)$ that cannot be part of any cycle in $H$, contradicting that $H$ is 2EC. If $H_{i}^{\prime}(v)$ is also a single node, then $H_{i}$ is of type $\mathsf{B2}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}(v)$ on the simple path between $C_{i}(u)$ and $C_{i}(v)$. If $H_{i}^{\prime}(v)$ contains an edge $e\in E(H_{i}^{\prime}(v))\setminus F$, then $e$ cannot belong to any cycle of $H$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C2}$.
Finally, assume that $H_{i}$ is composed of three connected components. Let $H_{i}^{\prime}(u)$, $H_{i}^{\prime}(v)$, and $H_{i}^{\prime}(w)$ denote the trees of $H_{i}^{\prime}$ that contain $u$, $v$, and $w$, respectively. Using the same argument as before, we can show that each tree is composed of a single node, and, thus, $H_{i}$ is of type $\mathsf{C3}$. ∎
@@
In Algorithm 2 it holds that $\mathsf{opt}_{1}^{\min}\geq 8$.
@@
Consider any feasible 2-ECSS solution $H_{1}$ for $G_{1}$, and let $H^{\prime}_{1}=H_{1}\setminus\{u,v,w\}$ be the corresponding solution for $G_{1}|\{u,v,w,\}$. Since $G_{1}|\{u,v,w,\}$ contains at least $8$ vertices and $H^{\prime}_{1}$ is a feasible 2-ECSS solution, $H^{\prime}_{1}$ must contain at least $8$ edges. Thus $|H_{1}|\geq|H^{\prime}_{1}|\geq 8$. ∎
@@
If Algorithm 2 reaches Algorithm 2 and $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, then $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1\geq\mathsf{opt}_{1}^{\min}+3$.
Proof.
If $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists and $\mathsf{opt}_{1}^{\mathsf{A}}\leq\alpha\cdot\mathsf{opt}_{1}^{\min}$ then $\mathsf{OPT}_{1}^{\mathsf{A}}$ is 2EC and $\alpha$-contractible. However, since we assume that we reached Algorithm 2, we have that $|V_{1}|\leq\frac{2}{\varepsilon}-4\leq\frac{4}{\varepsilon}$. Thus, this is a contradiction to Proposition 19, as we would have contracted $G_{1}$ in Algorithm 1 earlier. Thus, it must be that $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1$. Then, $\alpha\cdot\mathsf{opt}_{1}^{\min}\geq\mathsf{opt}_{1}^{\min}+2$, since $\alpha\geq\frac{5}{4}$ and $\mathsf{opt}_{1}^{\min}\geq 8$ by Lemma 21. ∎
We have an immediate corollary from this lemma, which justifies that we do not need to consider the case $t^{\min}=\mathsf{A}$ in Algorithm 2.
Corollary 23.
If Algorithm 2 reaches Algorithm 2, then $t^{\min}\neq\mathsf{A}$.
Lemma 24.
Let $G=(V,E)$ be a (multi-)graph that does not contain cut nodes or non-isolating 2-vertex cuts, and let $\{u,v,w\}$ be a large 3-vertex cut of $G$. Let $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$, and let $G_{1}\coloneq G[V_{i}\cup\{u,v,w\}]$ and $G_{2}\coloneq G[V_{i}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$. For $i\in\{1,2\}$, let $H_{i}\subseteq E(G_{i})$, let $H_{i}(z)$ be the connected component of $H_{i}$ that contains $z\in V(H_{i})$ and let $H_{i}$ be such that $V(H_{i})=V(H_{i}(u))\cup V(H_{i}(v))\cup V(H_{i}(w))$, i.e., each vertex $y\in V(H_{i})$ is connected to $u$, $v$, or $w$ in $H_{i}$. Then, the following holds.
@@
We prove the first claim, and note that the second claim follows from applying the first claim twice. Let $i\in\{1,2\}$, $x\in\{u,v,w\}$ and $H_{i}(x)\neq H_{i}$. If $H_{i}(x)$ is composed of the single vertex $x$ such that there is no edge between $x$ and $H_{i}\setminus H_{i}(x)$, then $\{u,v,w\}\setminus\{x\}$ must be a non-isolating two-vertex cut in $G$, a contradiction. Otherwise, that is, $H_{i}(x)$ consists of at least two vertices, and there is no edge between $H_{i}(x)$ and $H_{i}\setminus H_{i}(x)$ then $x$ is a cut node in $G$ (separating $V(H_{i}(x))\setminus\{x\}$ from $V\setminus V(H_{i}(x))$), a contradiction. ∎

### v2

@@ A.2.1 Auxiliary Lemmas for Algorithm˜2
If Algorithm˜1 executes Algorithm˜2, then $G$ contains no self loops, parallel edges, irrelevant edges, or $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices.
Lemma 23 (Lemma A.4 in [GGA23]).
Let $G$ be the graph when Algorithm˜1 executes Algorithm˜2, $\{u,v\}$ be a non-isolating 2-vertex cut of $G$, $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$, and $H$ be a 2EC spanning subgraph of $G$. Let $G_{i}=G[V_{i}\cup\{u,v\}]$ and $H_{i}=E(G_{i})\cap H$ for $i\in\{1,2\}$. The following statements are true.
  1. 1.
    Both $H_{1}$ and $H_{2}$ are of a type in $\mathcal{T}_{2}$ with respect to $\{u,v\}$. Furthermore, if one of the two is of type $\mathsf{C}$, then the other must be of type $\mathsf{A}$.
  2. 2.
    If $H_{i}$, for an $i\in\{1,2\}$, is of type $\mathsf{C}$, then there exists one edge $f\in E(G_{i})$ such that $H_{i}^{\prime}\coloneq H_{i}\cup\{f\}$ is of type $\mathsf{B}$. As a consequence, there exists a 2EC spanning subgraph $H^{\prime}$ where $H_{i}^{\prime}\coloneq H^{\prime}\cap E(G_{i})$ is of type $\mathsf{A}$ or $\mathsf{B}$.
Proof.
First note that $uv\notin E(G)$ since there are no irrelevant edges by Proposition˜22. We first prove the first claim. We prove it for $H_{1}$, the other case being symmetric. First notice that if $H_{1}$ contains a connected component not containing $u$ nor $v$, $H$ would be disconnected. Hence, $H_{1}$ consists of one connected component or two connected components $H_{1}(u)$ and $H_{1}(v)$ containing $u$ and $v$, respectively.
Suppose first that $H_{1}$ consists of one connected component. Let us contract the 2EC components of $H_{1}$, hence obtaining a tree $T$. If $T$ consists of a single vertex, $H_{1}$ is of type $\mathsf{A}$. Otherwise, consider the path $P$ (possibly of length 0) between the super nodes resulting from the contraction of the 2EC components $C(u)$ and $C(v)$ containing $u$ and $v$, respectively. Assume by contradiction that $T$ contains an edge $e$ not in $P$. Then $e$ does not belong to any cycle of $H$, contradicting the fact that $H$ is 2EC. Thus, $H_{1}$ is of type $\mathsf{B}$ (in particular $P$ has length at least 1 since $H_{1}$ is not 2EC).
Assume now that $H_{1}$ consists of 2 connected components $H_{1}(u)$ and $H_{1}(v)$. Let $T_{u}$ and $T_{v}$ be the two trees obtained by contracting the 2EC components of $H_{1}(u)$ and $H_{1}(v)$, respectively. By the same argument as before, the 2-edge connectivity of $H$ implies that these two trees contain no edge. Hence, $H_{1}$ is of type $\mathsf{C}$.
The second part of the first claim follows easily since if $H_{1}$ is of type $\mathsf{C}$ and $H_{2}$ is not of type $\mathsf{A}$, then $H$ would either be disconnected or it would contain at least one bridge edge, because $uv\notin E(G)$.
We now move to the second claim of the lemma. There must exist an edge $f$ in $G_{i}$ between the two connected components of $H_{i}$ since otherwise at least one of $u$ or $v$ would be a cut vertex. Clearly, $H^{\prime}_{i}\coloneq H_{i}\cup\{f\}$ satisfies the claim. For the second part of the claim, consider any 2EC spanning subgraph $H^{\prime}$ and let $H^{\prime}_{i}\coloneq H^{\prime}\cap E(G_{i})$. The claim holds if there exists one such $H^{\prime}$ where $H^{\prime}_{i}$ is of type $\mathsf{A}$ or $\mathsf{B}$, hence assume that this is not the case. Hence, $H^{\prime}_{i}$ is of type $\mathsf{C}$ by the first part of this lemma. The same argument as above implies the existence of an edge $f$ in $G_{i}$ such that $H^{\prime\prime}_{i}\coloneq H^{\prime}_{i}\cup\{f\}$ is of type $\mathsf{B}$, implying the claim for $H^{\prime\prime}\coloneq H^{\prime}\cup\{f\}$. ∎
@@
In Algorithm˜2 it holds that $\mathsf{opt}_{1}^{\min}\geq 3$.
Proof.
Consider any feasible 2-ECSS solution $H_{1}$ for $G_{1}$, and let $H^{\prime}_{1}=H_{1}\setminus\{u,v\}$ be the corresponding solution for $G_{1}|\{u,v\}$. Since $G_{1}|\{u,v\}$ contains at least $3$ vertices and $H^{\prime}_{1}$ is a feasible 2-ECSS solution, $H^{\prime}_{1}$ must contain at least $3$ edges. Thus, $|H_{1}|\geq|H^{\prime}_{1}|\geq 3$. ∎
We have an immediate corollary from Lemma˜21 and $\alpha>1$.
Corollary 25.
If Algorithm˜2 reaches Algorithm˜2 and $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, then $\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\min}+2$. In particular, $t^{\min}\neq\mathsf{A}$.
Lemma 26 (Claim 1 in [GGA23]).
If Algorithm˜2 is executed on instance $G$, there exists an optimal solution $\mathsf{opt}(G)$ such that $\mathsf{OPT}_{1}$ is of type $\mathsf{B}$ or $\mathsf{C}$.
Proof.
For the sake of contradiction, consider any optimal solution $\mathsf{OPT}(G)$ and assume that $\mathsf{OPT}_{1}$ is of type $\mathsf{A}$. Lemma˜23 implies that every feasible solution must use at least $\mathsf{opt}_{1}^{\min}$ edges from $G_{1}$. Moreover, $\mathsf{opt}_{1}^{\mathsf{A}}>\alpha\cdot\min\{\mathsf{opt}_{1}^{\mathsf{B}},\mathsf{opt}_{1}^{\mathsf{C}}\}$ by Lemma˜21. If $\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{B}}+1$, then we obtain an alternative optimum solution with the desired property by taking $\mathsf{OPT}_{1}^{\mathsf{B}}\cup\mathsf{OPT}_{2}$, and, in case $\mathsf{OPT}_{2}$ is of type $\mathsf{C}$, by adding one edge $f$ between the connected components of $\mathsf{OPT}_{2}$ whose existence is guaranteed by Lemma˜23. Otherwise, that is, $\mathsf{opt}_{1}^{\mathsf{A}}\leq\mathsf{opt}_{1}^{\mathsf{B}}$, we must have $\mathsf{opt}_{1}^{\mathsf{A}}>\alpha\cdot\mathsf{opt}_{1}^{\mathsf{C}}$. Specifically, $\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C}}+1\geq 4$ by Lemma˜24. Lemma˜23 also implies $\mathsf{opt}_{1}^{\mathsf{B}}\leq\mathsf{opt}_{1}^{\mathsf{C}}+1$, which together gives $\mathsf{opt}_{1}^{\mathsf{A}}=\mathsf{opt}_{1}^{\mathsf{B}}=\mathsf{opt}_{1}^{\mathsf{C}}+1$. Then, $\mathsf{opt}_{1}^{\mathsf{A}}>\alpha\cdot\mathsf{opt}_{1}^{\mathsf{C}}=\alpha(\mathsf{opt}_{1}^{\mathsf{A}}-1)$ implies $\mathsf{opt}_{1}^{\mathsf{A}}\leq 5$ since $\alpha\geq\frac{6}{5}$. We established that $\mathsf{opt}_{1}^{\mathsf{A}}\in\{4,5\}$.
If $\mathsf{opt}_{1}^{\mathsf{A}}=4$, $\mathsf{OPT}_{1}^{\mathsf{A}}$ has to be a $4$-cycle C=uavbuC=u-a-v-b-u. Notice that the edge $ab$ cannot exist since otherwise $\mathsf{opt}_{1}^{\mathsf{B}}=3$ due to the path uabvu-a-b-v. Then every feasible solution must include the 4 edges of $C$ to guarantee degree at least 2 on $a$ and $b$. But this makes $C$ an $\alpha$-contractible subgraph on $4\leq\frac{4}{\varepsilon}$ vertices, which is a contradiction to Proposition˜22.
If $\mathsf{opt}_{1}^{\mathsf{A}}=5$, the minimality of $\mathsf{OPT}_{1}^{\mathsf{A}}$ implies that $\mathsf{OPT}_{1}^{\mathsf{A}}$ is a $5$-cycle, say C=uavbcuC=u-a-v-b-c-u. Observe that the edge $ab$ cannot exist, since otherwise $\{va,ab,bc,cu\}$ would be a type $\mathsf{B}$ solution in $G_{1}$ of size $4$, contradicting $\mathsf{opt}_{1}^{\mathsf{B}}=\mathsf{opt}_{1}^{\mathsf{B}}=5$. A symmetric construction shows that edge $ac$ cannot exist. Hence, every feasible solution restricted to $G_{1}$ must include the edges $\{au,av\}$ and furthermore at least $3$ more edge incident to $b$ and $c$, so at least $5$ edges altogether. This implies that $C$ is an $\alpha$-constractible subgraph of size $5\leq\frac{4}{\varepsilon}$ vertices, which is a contradiction to Proposition˜22. ∎
Lemma 27 (Part of the proof of Lemma A.6 in [GGA23]).
If the condition in Algorithm˜2 in Algorithm˜2 applies and both, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are 2EC spanning subgraphs of $G^{\prime}_{1}$ and $G^{\prime}_{2}$, respectively, then there exist edges $F^{\prime}$ with $|F^{\prime}|\leq 2$ such that $H^{\prime}$ is a 2EC spanning subgraph of $G$.
Proof.
First note that $H^{\prime}_{i}$, $i\in\{1,2\}$, must be of type $\mathsf{A}$, $\mathsf{B}$, or $\mathsf{C}$. If one of them is of type $\mathsf{A}$ or they are both of type $\mathsf{B}$, then $H^{\prime}_{1}\cup H^{\prime}_{2}$ is 2EC. If one of them is of type $\mathsf{B}$, say $H^{\prime}_{1}$, and the other of type $\mathsf{C}$, say $H^{\prime}_{2}$, then let $H^{\prime}_{2}(u)$ and $H^{\prime}_{2}(v)$ be the two 2EC components of $H^{\prime}_{2}$ containing $u$ and $v$, respectively. By Lemma˜23, there must exist an edge $f\in E(G_{2})$ between $H^{\prime}_{2}(u)$ and $H^{\prime}_{2}(v)$. Hence, $H^{\prime}_{1}(u)\cup H^{\prime}_{2}(v)\cup\{f\}$ is 2EC. The remaining case is that $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are both of type $\mathsf{C}$. In this case, $H^{\prime}_{1}\cup H^{\prime}_{2}$ consists of precisely two 2EC components $H^{\prime}(u)$ and $H^{\prime}(v)$ containing $u$ and $v$, respectively. Since $G$ is 2EC, there must exist two edges $f$ and $g$ between $H^{\prime}(u)$ and $H^{\prime}(v)$ such that $H^{\prime}_{1}\cup H^{\prime}_{2}\cup\{f,g\}$ is 2EC. Thus, in all cases a desired set $F^{\mathsf{C}}$ of size at most 2 exists such that $H^{\prime}$ is a 2-ECSS of $G$. ∎
Lemma 28 (Part of the proof of Lemma A.6 in [GGA23]).
If the condition in Algorithm˜2 in Algorithm˜2 applies and $\mathsf{RED}(G_{2}^{\mathsf{B}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B}}$, then $H^{\mathsf{B}}$ is a 2EC spanning subgraph of $G$.
Proof.
Note that $\mathsf{RED}(G_{2}^{\mathsf{B}})$ must contain the two dummy edges $wv$ and $wu$, which induce a type $\mathsf{B}$ graph. Hence, by Lemma˜23, $H^{\mathsf{B}}_{2}$ is of type $\mathsf{A}$ or $\mathsf{B}$. Thus, $H^{\mathsf{B}}=\mathsf{OPT}_{1}^{\mathsf{B}}\cup H^{\mathsf{B}}$ is a 2-ECSS of $G$. ∎
Lemma 29 (Part of the proof of Lemma A.6 in [GGA23]).
If the condition in Algorithm˜2 in Algorithm˜2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C}}$, then there exists an edge set $F^{\mathsf{C}}$ with $|F^{\mathsf{C}}|\leq 1$ such that $H^{\mathsf{C}}$ is a 2EC spanning subgraph of $G$.
Proof.
Note that $H_{2}^{\mathsf{C}}$ must be of type $\mathsf{A}$ or $\mathsf{B}$. Lemma˜23 guarantees the existence of an edge $f\in E(G_{2})$ such that $\mathsf{OPT}_{1}^{\mathsf{C}}\cup\{f\}$ is of type $\mathsf{B}$ for $G_{1}$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C}}\cup H_{2}^{\mathsf{C}}\cup\{f\}$ is a 2-ECSS of $G$. ∎
Finally, we prove that the reductions of Algorithm˜2 preserve the approximation factor.
Lemma 30 (Part of the proof of Lemma A.7 in [GGA23]).
Let $G$ be a 2-ECSS instance. If every recursive call to $\mathsf{RED}(G^{\prime})$ in Algorithm˜2 on input $G$ satisfies $\mathsf{red}(G^{\prime})\leq\alpha\cdot\mathsf{opt}(G^{\prime})+4\varepsilon\cdot|V(G^{\prime})|-4$, then it holds that $\mathsf{red}(G)\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\cdot|V(G)|-4$.
Proof.
Let $\mathsf{OPT}_{i}\coloneq\mathsf{OPT}(G)\cap E(G_{i})$ and $\mathsf{opt}_{i}\coloneq\lvert\mathsf{OPT}_{i}\rvert$ for $i\in\{1,2\}$. Note that $\mathsf{opt}(G)=\mathsf{opt}_{1}+\mathsf{opt}_{2}$. For the remainder of this proof, all line references (if not specified differently) refer to Algorithm˜2. We use in every case that $\mathsf{opt}_{1}\geq 2$ (cf. Lemma˜24). We distinguish in the following which reduction the algorithm uses.
(Algorithm 2: $|V_{1}|>\frac{2}{\varepsilon}-4$)
Note that for $i\in\{1,2\}$ we have that $\mathsf{OPT}_{i}$ is a feasible solution for $G_{i}^{\prime}$ and therefore $\mathsf{opt}_{i}\geq\mathsf{opt}(G_{i}^{\prime})$. Using the assumption of the lemma, Lemma˜27 and $\lvert V(G_{1}^{\prime})\rvert+\lvert V(G_{2}^{\prime})\rvert=\lvert V(G)\rvert$, we conclude that
$\mathsf{red}(G)=\lvert H^{\prime}_{1}\rvert+\lvert H^{\prime}_{2}\rvert+\lvert F^{\prime}\rvert$ $\leq\alpha\cdot(\mathsf{opt}_{1}+\mathsf{opt}_{2})+4\varepsilon(\lvert V(G_{1}^{\prime})\rvert+\lvert V(G_{2}^{\prime})\rvert)-8+2$
$\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\cdot(\lvert V(G)\rvert)-4\ .$
(Algorithm 2: $t^{\min}=\mathsf{B}$)
Note that $\mathsf{opt}_{2}(G_{2}^{\mathsf{B}})\leq\mathsf{opt}_{2}+2$, because we can turn $\mathsf{OPT}_{2}$, which we can assume to be of type $\mathsf{A}$ or $\mathsf{B}$ by Lemma˜26 and Lemma˜23, using both dummy edges $\{uw,vw\}$ into a solution for $G_{2}^{\mathsf{B}}$. Thus,
$\mathsf{red}(G)=\lvert H^{\mathsf{B}}\rvert$ $\leq\mathsf{opt}_{1}^{\mathsf{B}}+\lvert H^{\mathsf{B}}\rvert\leq\mathsf{opt}_{1}+(\alpha\cdot\mathsf{opt}_{2}(G_{2}^{\mathsf{B}})+4\varepsilon|V(G)|-4)-2$
$\leq\mathsf{opt}_{1}+(\alpha\cdot(\mathsf{opt}_{2}+2)+4\varepsilon|V(G)|-4)-2$
$\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4+(\alpha-1)(2-\mathsf{opt}_{1})\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4\ .$
(Algorithm 2: $t^{\min}=\mathsf{C}$)
We distinguish the following cases depending on which type combination of Lemma˜23 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$. We exclude the case that $\mathsf{OPT}_{1}$ is of type $\mathsf{A}$ due to Lemma˜26. Moreover, Lemma˜29 gives that $|F^{\mathsf{C}}|\leq 1$.
  1. (a)
    $(\mathsf{B},\{\mathsf{A},\mathsf{B}\})$. In this case, $\mathsf{opt}_{1}^{\mathsf{B}}$ exists but $t^{\min}=\mathsf{C}$. Thus, the definition of Algorithm˜2 and Definition˜14 imply that $\mathsf{opt}_{1}^{\mathsf{C}}\leq\mathsf{opt}^{\mathsf{B}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}_{2}(G_{2}^{\mathsf{C}})\leq\mathsf{opt}_{2}+1$, because we can turn $\mathsf{OPT}_{2}$ with the dummy edge $\{uv\}$ into a solution for $G_{2}^{\mathsf{C}}$. Thus,
    $\mathsf{red}(G)=\lvert H^{\mathsf{C}}\rvert$ $\leq\mathsf{opt}_{1}^{\mathsf{C}}+\lvert H^{\mathsf{C}}\rvert+\lvert F^{\mathsf{C}}\rvert$
    $\leq(\mathsf{opt}_{1}-1)+(\alpha\cdot(\mathsf{opt}_{2}+1)+4\varepsilon|V(G)|-4)-1+1$
    $\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4+(\alpha-1)(1-\mathsf{opt}_{1})\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4\ .$
  2. (b)
    $(\mathsf{C},\mathsf{A})$. In this case, $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{C}}$ and $\mathsf{OPT}_{2}$ is a feasible solution for $G_{2}^{\mathsf{C}}$. Thus,
    $\mathsf{red}(G)=\lvert H^{\mathsf{C}}\rvert$ $\leq\mathsf{opt}_{1}^{\mathsf{C}}+\lvert H^{\mathsf{C}}\rvert+\lvert F^{\mathsf{C}}\rvert$
    $\leq\mathsf{opt}_{1}+(\alpha\cdot\mathsf{opt}_{2}+4\varepsilon|V(G)|-4)-1+1\leq\alpha\mathsf{opt}(G)+4\varepsilon|V(G)|-4\ .$
This completes the proof of the lemma. ∎
A.2.2 Auxiliary Lemmas for Algorithm˜3
Proposition 31.
If Algorithm˜1 executes Algorithm˜3, then $G$ contains no non-isolating 2-vertex cuts, self loops, parallel edges, irrelevant edges, or $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices.
Lemma 32.
@@ A.2.2 Auxiliary Lemmas for Algorithm˜3
Let $G$ be the graph when Algorithm˜1 executes Algorithm˜3, $\{u,v,w\}$ be a large 3-vertex cut of $G$, $(V_{1},V_{2})$ be partition of $V\setminus\{u,v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$, and $H$ be a 2EC spanning subgraph of $G$. Let $G_{i}=G[V_{i}\cup\{u,v,w\}]$ and $H_{i}=E(G_{i})\cap H$ for $i\in\{1,2\}$. Then, $H_{1}$ and $H_{2}$ are of a type in $\mathcal{T}_{3}$ with respect to $\{u,v,w\}$ and $G_{1}$ and $G_{2}$, respectively.
@@
First, assume that $H_{i}$ is a single connected component. In this case $H_{i}^{\prime}$ is a tree. If $H_{i}^{\prime}$ is composed of a single vertex, $H_{i}$ is of type $\mathsf{A}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}$ that are on any simple path between $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$. If $H_{i}^{\prime}$ contains an edge $e\in E(H_{i}^{\prime})\setminus F$, then $e$ cannot belong to any cycle of $H$, because it does not belong to a cycle in $H_{i}$ and is not on a simple path between any two vertices of the cut $\{u,v,w\}$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C1}$ if $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$ are distinct, and of type $\mathsf{B1}$ otherwise.
Second, assume that $H_{i}$ is composed of two connected components. Let $H_{i}^{\prime}(v)$ and $H_{i}^{\prime}(w)$ denote the two trees in $H_{i}^{\prime}$ such that w.l.o.g. $u$ and $v$ are in $H_{i}^{\prime}(v)$ and $w$ is in $H_{i}^{\prime}(w)$. First note that $H_{i}^{\prime}(w)$ must be a single vertex. Otherwise, there is an edge $e$ in $H_{i}^{\prime}(w)$ that cannot be part of any cycle in $H$, contradicting that $H$ is 2EC. If $H_{i}^{\prime}(v)$ is also a single vertex, then $H_{i}$ is of type $\mathsf{B2}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}(v)$ on the simple path between $C_{i}(u)$ and $C_{i}(v)$. If $H_{i}^{\prime}(v)$ contains an edge $e\in E(H_{i}^{\prime}(v))\setminus F$, then $e$ cannot belong to any cycle of $H$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C2}$.
Finally, assume that $H_{i}$ is composed of three connected components. Let $H_{i}^{\prime}(u)$, $H_{i}^{\prime}(v)$, and $H_{i}^{\prime}(w)$ denote the trees of $H_{i}^{\prime}$ that contain $u$, $v$, and $w$, respectively. Using the same argument as before, we can show that each tree is composed of a single vertex, and, thus, $H_{i}$ is of type $\mathsf{C3}$. ∎
@@
In Algorithm˜3 it holds that $\mathsf{opt}_{1}^{\min}\geq 8$.
@@
Consider any feasible 2-ECSS solution $H_{1}$ for $G_{1}$, and let $H^{\prime}_{1}=H_{1}\setminus\{u,v,w\}$ be the corresponding solution for $G_{1}|\{u,v,w\}$. Since $G_{1}|\{u,v,w\}$ contains at least $8$ vertices and $H^{\prime}_{1}$ is a feasible 2-ECSS solution, $H^{\prime}_{1}$ must contain at least $8$ edges. Thus, $|H_{1}|\geq|H^{\prime}_{1}|\geq 8$. ∎
We have an immediate corollary from the above lemma, Lemma˜21, and $\alpha\geq\frac{5}{4}$.
Corollary 34.
If Algorithm˜3 reaches Algorithm˜3 and $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, then $\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\min}+3$. In particular, $t^{\min}\neq\mathsf{A}$.
@@
Let $G=(V,E)$ be a (multi-)graph that does not contain cut vertices or non-isolating 2-vertex cuts, and let $\{u,v,w\}$ be a large 3-vertex cut of $G$. Let $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$, and let $G_{1}\coloneq G[V_{i}\cup\{u,v,w\}]$ and $G_{2}\coloneq G[V_{i}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$. For $i\in\{1,2\}$, let $H_{i}\subseteq E(G_{i})$, let $H_{i}(z)$ be the connected component of $H_{i}$ that contains $z\in V(H_{i})$ and let $H_{i}$ be such that $V(H_{i})=V(H_{i}(u))\cup V(H_{i}(v))\cup V(H_{i}(w))$, i.e., each vertex $y\in V(H_{i})$ is connected to $u$, $v$, or $w$ in $H_{i}$. Then, the following holds.
@@
We prove the first claim, and note that the second claim follows from applying the first claim twice. Let $i\in\{1,2\}$, $x\in\{u,v,w\}$ and $H_{i}(x)\neq H_{i}$. If $H_{i}(x)$ is composed of the single vertex $x$ such that there is no edge between $x$ and $H_{i}\setminus H_{i}(x)$, then $\{u,v,w\}\setminus\{x\}$ must be a non-isolating two-vertex cut in $G$, a contradiction. Otherwise, that is, $H_{i}(x)$ consists of at least two vertices, and there is no edge between $H_{i}(x)$ and $H_{i}\setminus H_{i}(x)$ then $x$ is a cut vertex in $G$ (separating $V(H_{i}(x))\setminus\{x\}$ from $V\setminus V(H_{i}(x))$), a contradiction. ∎